# PropMolFlow: Property-guided Molecule Generation with Geometry-Complete Flow Matching

### v3

@@ Methods
Joint Flow Matching for Molecule Generation
Ideally, new molecules would be generated by sampling directly from the probability distribution over valid molecular graphs $q(x)$. However, this target distribution lacks a tractable analytical form and can only be approximated through a generative process. To this end, we adopt flow matching to learn an approximation of $q(x)$. This flow matching is parameterized by an SE(3) equivariant graph neural network built on top of FlowMol architecture [26], details of which are provided in Supplemental Section LABEL:sec:appendix_model. In flow matching, the objective is to learn a time-dependent velocity field $u_{t}^{\theta}$ parameterized by a neural network $\theta$, which defines a probability path $p_{t}(x):[0,1]\times x_{0}\rightarrow x_{1}$ from a base distribution, $p_{0}(x)$ to a target distribution, $p_{1}(x)=q(x)$.iiiWe use subscripts to denote time dependence, e.g., $p_{t}(x)$ The flow induced by the velocity field is given by the ODE $\dot{x}=u_{t}\left(x\right)$, which describes the evolution of samples along the probability paths. Designing a flow matching process thus involves two key steps: 1) specifying a probability path that satisfies the boundary conditions $x_{0}\sim p_{0}$ and $x_{1}\sim p_{1}$, and 2) training a neural network $\theta$ to approximate the velocity field that transports samples along this path.
In practice, while accurately modeling the exact velocity field is challenging, conditioning it on a random variable $z$ leads to a more tractable formulation. Importantly, optimizing the conditional velocity field is equivalent to optimizing the marginal velocity field [40]. The endpoint reformulation of [24] reparametrizes the flow matching problem by focusing on learning a denoising function rather than the full time-dependent velocity field $u_{t}^{\theta}$. Specifically, the method learns a denoising function $x^{\theta}_{1|t}$ that predicts the final state $x_{1}$ from the intermediate state $x_{t}$. The training objective then minimizes the discrepancy between the predicted final sample and the true final sample:
@@
Discrete graph features are handled via CTMC as developed by Campell et al. [20] and Gat et al. [41]. In this setting, Eq. (3) is reformulated to minimize a cross-entropy loss over discrete variables:
@@
where $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=0$. The functions $\alpha_{t}$ and $\beta_{t}$, known as the interpolant schedule, control the rate of mixing between the base and target distributions. A typical and effective choice of interpolant schedule for both molecule and materials generation is a linear schedule: $\alpha_{t}=(1-t)$ and $\beta_{t}=t$ [26, 18]. Other variants of interpolants are discussed in detail in [17], and they have been validated in the context of inorganic crystals generation in our OMatG framework [18]. The base distribution for the atomic position flow is a standard Gaussian distribution $p_{0}(X)=\prod_{i=1}^{N}\mathcal{N}(X_{0}^{i}|\mathbf{0},\mathbb{I}_{3})$. A denoising network is then trained to produce the final atomic positions $X^{\theta}_{1|t}$, minimizing the following loss:
@@
Atom types, charge and bond orders are modeled using CTMC-based flows [20]. For each categorical variable, a ‘mask’ token is added as an additional discrete state. For example, consider atomic identities $A=\{A^{1},A^{2},\ldots,A^{N}\}$ where each $A^{i}$ takes values in the set $\{1,2,\ldots,n_{A},M\}$. Here, $n_{A}$ is the number of predefined atom types and $M$ is the mask token. The joint probability path over all atomic identities is factorized into independent contributions from each atom. The conditional probability path for atom $i$ is defined as:
@@
where $\alpha_{t}$ and $\beta_{t}$ are the same interpolant schedules defined in Eq. (6). A linear interpolant is effective for categorical variables as well [26, 18]. Here, $\delta(i,j)$ denotes the Kronecker delta, which is 1 if $i=j$ and 0 otherwise. The base distribution corresponds to all atoms being initialized in the masked state. This formulation implies that at time $t$, the atom identity $A^{i}_{t}$ has a probability $\alpha_{t}$ of being in its final state $A_{1}^{i}$ and a probability $\beta_{t}$ of being in the masked state $M$.
@@
$p^{i}_{t+\Delta t}(j|A^{i}_{t})$ =$=\left\{\begin{array}[]{lcl}R^{i}(A^{i}_{t},j)&\mbox{for}&j\neq A_{t}^{i}\\ 1+R^{i}(A^{i}_{t},A_{t}^{i})&\mbox{for}&j=A_{t}^{i}\end{array}\right.$ (11)
@@
where $R(A^{i}_{t},j)$ is the rate matrix specifying the transition probability from atom type $A^{i}_{t}$ to $j$ [20]. To ensure proper normalization of the transition probabilities at time $t+\Delta t$, the rate matrix must satisfy $R^{i}(A^{i}_{t},A^{i}_{t})=-\sum_{j\neq A^{i}_{t}}R(A^{i}_{t},j)$ so that the total probability mass remains normalized.
An additional stochastic term $\eta$ can be introduced into the rate matrix to form the modified rate matrix $R_{t}^{\eta}:=R_{t}^{*}+\eta R_{t}^{\mathrm{DB}}$. This new rate matrix generates the same marginal probability path if $R_{t}^{\mathrm{DB}}$ satisfies detailed balance over the conditional probability path $p_{t|1}$. By including the stochasticity term and choosing a linear interpolant ($\alpha_{t}=1-t$ and $\beta_{t}=t$), the rate matrix for atom $i$ becomes:
@@
This formulation implies that if atom $i$ is in the masked state at time $t$, the probability of transitioning to atom type $j$ at time step $t+\Delta t$ is $\Delta t\frac{1+\eta t}{1-t}$. Conversely, if atom $i$ is in an unmasked state, the probability of transitioning back to a masked state is $\eta\Delta t$. In Eq. (13), the rate matrix can diverge as $t\to 1$. To address this issue, we employ a time-dependent loss function that avoids division-by-zero [42].
Additionally, low-temperature sampling—which rescales the prediction logits to sharpen the output distribution—is found to be critical for improving model performance. Specifically, the denoiser network’s output $p^{\theta}_{1|t}(A_{1}^{i}|A_{t})$ is transformed as follows:
@@
where $\nu$ is a hyperparameter. The modified probabilities $\widehat{p}^{\theta}_{1|t}(A_{1}^{i}|A_{t})$ are then employed in the discrete loss function defined in Eq. (4). When training PropMolFlow, we set $\eta=10$ and $\nu=0.05$. During inference, we observe that turning off the stochastic term leads to better performance. To ensure that the generated molecules satisfy the required symmetry constraints, the joint flow matching process is constructed with an invariant base distribution and an equivariant transition probability path. A formal derivation of the equivariance properties of the flow matching process is provided in Supplementary Section LABEL:sec:equivariantfm.
Property Embedding Methods
@@
Conditional generation can be framed as sampling a property-conditional target distribution $p_{1}(x|k)$ over molecular graphs, where $k$ is a property. Previous approaches to conditional generation concatenate a property value to the node features of a molecular graph, training the model jointly on the property and the graph [7, 8, 21]. As a result, the learned denoising process is conditioned on the property. This approach requires modifying the input dimension of the denoiser to include the additional property value, which is subsequently discarded in the final output. However, directly appending the property to the node features may oversimplify the interaction between property values and molecular graphs. As shown in Figure 1, our implementation maps a scalar property, $k$, to a high-dimensional property embedding space $P=\phi_{\mathrm{prop}}(k)$ using a MLP, $\phi_{\mathrm{prop}}$. The dimensionality of the property embedding matches that of the node scalar features, allowing seamless integration without requiring additional dimensional matching. To encode interactions between the property embedding and node features, we propose five distinct operations. Depending on the operation, if the combined feature vector $[A,C]$ maintains its original dimensionality after interacting with the property embedding, no further processing is applied; otherwise, an MLP is used to project it back to the original dimensionality. Since the property embedding interacts exclusively with node scalar features, leaving position vectors unchanged, the resulting property-conditioned graph retains SE(3) equivariance. Moreover, the loss function preserves its original form, and the trained SE(3) GNN defines a valid conditional generative process.
@@
In the ‘Sum’ operation, the property embedding is simply added to the node scalar features. The ‘Multiply’ operation performs an element-wise Hadamard product $\odot$’. To ensure the multiplicative factors remain within a bounded range, the property embedding is passed through a sigmoid function and then shifted by 0.5, resulting in multipliers in the interval $[0.5,1.5]$. Both ‘Sum and Multiply’ preserve the original feature dimensionality. In contrast, operations involving ‘Concatenate’ increase the dimensionality of the node scalar features, and an MLP transformation $\varphi_{\theta}$ is required to project the combined features back to the original dimension. For the composite operations, ‘Concatenate + Sum’ and ‘Concatenate + Multiply’, the property embedding is first concatenated to the node features, followed by the respective arithmetic operation. In both cases, the final output has the desired dimensionality, either by design or via the subsequent MLP transformation.
Gaussian Expansion for Augmented Property Embedding
@@
We incorporate a non-trainable Gaussian expansion layer to enrich the property representations prior to being mapped to a property embedding via an MLP $\phi_{\mathrm{GE}}$ [29]. The Gaussian expansion maps a scalar property value into a fixed-length vector using a set of Gaussian basis functions. Given a property $k$ with value $\tau_{k}$, the expanded representation takes the form:
@@
Figure 1d shows an example of a Gaussian expansion for the HOMO-LUMO gap. In this example, five Gaussian basis functions ($n_{g}=5$) are used, with centers evenly spaced between the minimum and maximum gap values in the QM9 dataset: $\tau_{\mathrm{min}}=0.0246$ and $\tau_{\mathrm{max}}=0.6221$ Hartree. The width of each Gaussian is set by the spacing parameter $d=(\tau_{\mathrm{max}}-\tau_{\mathrm{min}})/(n_{g}-1)$. The top panel of Figure 1d shows the five Gaussian basis functions, with their centers marked by vertical dashed gray lines. In the bottom panel, we pick three example property values, namely 0.1, 0.3, and 0.5 Hartree, and show the expanded feature vectors as bar plots indicating the activation of each Gaussian basis function. For instance, if the gap is 0.1 Hartree, the first two Gaussian functions (centered closest to 0.1) have higher responses than the remaining functions, whose centers are farther away. Similar activation patterns are observed for the other two property values.
Datasets and Baselines
Datasets.
@@
We used RDKit’s sanitization function [33] to filter chemically valid molecules on the QM9 SDF file with bond and charge fix, resulting in a curated set of 133k molecules. After random shuffling, we split the dataset into 100k training, 20k validation, and 13k test samples. For property-guided generation, we considered six molecule properties: polarizability ($\alpha$), HOMO-LUMO gap, HOMO energy, LUMO energy, dipole moment ($\mu$), and heat capacity ($C_{v}$). Precise definitions of each property can be found in Supplementary Section LABEL:sec:more_details. We did not evaluate PropMolFlow on the GEOM-Drugs dataset, which is commonly used for unconditional generation and includes larger molecules, because it only provides quantum-mechanical energy values and lacks other relevant properties [43]. PropMolFlow models and GVP property predictors are trained on disjoint 50k and 50k datasets from the 100k training data.
Baselines.

### v4

@@ Methods
Joint flow matching
Molecules are represented by fully-connected graphs $G$. Graph nodes encode the atomic types $A$, charges $C$ and positions $X$ of a molecule, and graph edges include bond orders $E$. Ideally, new molecules would be generated by sampling directly from the probability distribution over valid molecular graphs $q(x)$, where $x$ is an arbitrary sample. However, this target distribution lacks a tractable analytical form and can only be approximated through a generative process. To this end, we adopt flow matching to learn an approximation of $q(x)$. This flow matching is parameterized by an SE(3) equivariant graph neural network built on top of FlowMol architecture [dunn2024-2], details of which are provided in the subsequent ‘Model architecture’ section. In flow matching, the objective is to learn a time-dependent velocity field $u_{t}^{\theta}$ parameterized by a neural network $\theta$, which defines a probability path $p_{t}(x):[0,1]\times x_{0}\rightarrow x_{1}$ from a base distribution, $p_{0}(x)$ to a target distribution, $p_{1}(x)=q(x)$ (subscripts denote time dependence, for example, $p_{t}(x)$) The flow induced by the velocity field is given by the ODE $\dot{x}=u_{t}\left(x\right)$, which describes the evolution of samples along the probability paths. Designing a flow matching process thus involves two key steps: 1) specifying a probability path that satisfies the boundary conditions $x_{0}\sim p_{0}$ and $x_{1}\sim p_{1}$, and 2) training a neural network $\theta$ to approximate the velocity field that transports samples along this path.
In practice, while accurately modeling the exact velocity field is challenging, conditioning it on a random variable $z$ leads to a more tractable formulation. Importantly, optimizing the conditional velocity field is equivalent to optimizing the marginal velocity field [lipman2023]. The endpoint reformulation of [dunn2024-1] reparametrizes the flow matching problem by focusing on learning a denoising function rather than the full time-dependent velocity field $u_{t}^{\theta}$. Specifically, the method learns a denoising function $x^{\theta}_{1|t}$ that predicts the final state $x_{1}$ from the intermediate state $x_{t}$. The training objective then minimizes the discrepancy between the predicted final sample and the true final sample:
@@
Discrete graph features are handled via CTMC as developed by Campell et al. [campbell2024] and Gat et al. [gat2024]. In this setting, Eq. (1) is reformulated to minimize a cross-entropy loss over discrete variables:
@@
where $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=0$. The functions $\alpha_{t}$ and $\beta_{t}$, known as the interpolant schedule, control the rate of mixing between the base and target distributions. A typical and effective choice of interpolant schedule for both molecule and materials generation is a linear schedule: $\alpha_{t}=(1-t)$ and $\beta_{t}=t$ [dunn2024-2, hoellmer2025]. Other variants of interpolants are discussed in detail in [albergo2023, hoellmer2025]. The base distribution for the atomic position flow is a standard Gaussian distribution $p_{0}(X)=\prod_{i=1}^{N}\mathcal{N}(X_{0}^{i}|\mathbf{0},\mathbb{I}_{3})$. A denoising network is then trained to produce the final atomic positions $X^{\theta}_{1|t}$, minimizing the following loss:
@@
Atom types, charge and bond orders are modeled using CTMC-based discrete flow matching [campbell2024]. For each categorical variable, a ‘mask’ token is added as an additional discrete state. For example, consider atomic identities $A=\{A^{1},A^{2},\ldots,A^{N}\}$ where each $A^{i}$ takes values in the set $\{1,2,\ldots,n_{A},M\}$. Here, $n_{A}$ is the number of predefined atom types and $M$ is the mask token. The joint probability path over all atomic identities is factorized into independent contributions from each atom. The conditional probability path for atom $i$ is defined as:
@@
where $\alpha_{t}$ and $\beta_{t}$ are the same interpolant schedules defined in Eq. (4). A linear interpolant is effective for categorical variables as well [dunn2024-2, hoellmer2025]. Here, $\delta(i,j)$ denotes the Kronecker delta, which is 1 if $i=j$ and 0 otherwise. The base distribution corresponds to all atoms being initialized in the masked state. This formulation implies that at time $t$, the atom identity $A^{i}_{t}$ has a probability $\alpha_{t}$ of being in its final state $A_{1}^{i}$ and a probability $\beta_{t}$ of being in the masked state $M$.
@@
$p^{i}_{t+\Delta t}(j|A^{i}_{t})$ $=\left\{\begin{array}[]{lcl}R^{i}(A^{i}_{t},j)&\mbox{for}&j\neq A_{t}^{i}\\ 1+R^{i}(A^{i}_{t},A_{t}^{i})&\mbox{for}&j=A_{t}^{i}\end{array}\right$. (9)
@@
where $R(A^{i}_{t},j)$ is the rate matrix specifying the transition probability from atom type $A^{i}_{t}$ to $j$ [campbell2024]. To ensure proper normalization of the transition probabilities at time $t+\Delta t$, the rate matrix must satisfy $R^{i}(A^{i}_{t},A^{i}_{t})=-\sum_{j\neq A^{i}_{t}}R(A^{i}_{t},j)$ so that the total probability mass remains normalized.
An additional stochastic term $\eta$ can be introduced into the rate matrix to form the modified rate matrix $R_{t}^{\eta}:=R_{t}^{*}+\eta R_{t}^{\mathrm{DB}}$. This new rate matrix generates the same marginal probability path if $R_{t}^{\mathrm{DB}}$ satisfies detailed balance (‘DB’) over the conditional probability path $p_{t|1}$. By including the stochasticity term and choosing a linear interpolant ($\alpha_{t}=1-t$ and $\beta_{t}=t$), the rate matrix for atom $i$ becomes:
@@
This formulation implies that if atom $i$ is in the masked state at time $t$, the probability of transitioning to atom type $j$ at time step $t+\Delta t$ is $\Delta t\frac{1+\eta t}{1-t}$. Conversely, if atom $i$ is in an unmasked state, the probability of transitioning back to a masked state is $\eta\Delta t$. In Eq. (11), the rate matrix can diverge as $t\to 1$. To address this issue, we employ a time-dependent loss function that avoids division-by-zero [le2023].
Additionally, low-temperature sampling—which rescales the prediction logits to sharpen the output distribution—is found to be critical for improving model performance [dunn2024-2]. Specifically, the denoiser network’s output $p^{\theta}_{1|t}(A_{1}^{i}|A_{t})$ is transformed as follows:
@@
where $\nu$ is a hyperparameter. The modified probabilities $\widehat{p}^{\theta}_{1|t}(A_{1}^{i}|A_{t})$ are then employed in the discrete loss function defined in Eq. (2). When training PropMolFlow, we set $\eta=10$ and $\nu=0.05$. During inference, we observe that turning off the stochastic term leads to better performance. To ensure that the generated molecules satisfy the required symmetry constraints, the joint flow matching process is constructed with an invariant base distribution and an equivariant transition probability path parameterized by the SE(3)-GVP architecture. A self-contained formal derivation of probability invariance and equivariant generative process is provided in respective Supplementary Section LABEL:sec:equivariantfm and Supplementary Section LABEL:sec:se3_gvp.
Model architecture
We utilize the FlowMol architecture, which is implemented with PyTorch and Deep Graph Library [dunn2024-2]. Molecule updates are achieved through layers comprising Geometric Vector Perceptrons (GVP). Within each GVP, the molecule graph passes through a sequential steps of node feature update, node position update and edge feature update. Each node $i$ consists of a position $x_{i}\in\mathbb{R}^{3}$, scalar features $s_{i}\in\mathbb{R}^{d}$, and vector features $v_{i}\in\mathbb{R}^{c\times 3}$. The scalar feature is a concatenation of atom type and charge vectors; that is, $s_{i}:=[a_{i}:c_{i}]$ where ‘:’ defines a concatenation operation. Vector features are initialized at zeros, and despite not existing in the final outputs, their updates employ the vector cross-products that is not equivariant to reflections, hence SE(3) equivariant [dunn2024-1]. Each edge feature corresponds to the bond order and the permutation invariance of the scaler bond order is ensured by taking the sum of learned bond features from $i\to j$ and $j\to i$;i.e., $\widehat{e}^{ij}=\mathrm{MLP}(e_{ij}+e_{ji})$.
Node feature update.
The node feature uses graph convolution to update node scalar and vector features $s_{i},v_{i}$. It follows two steps: the first step generates scalar $m_{i\to j}^{(s)}$ and vector messages $m_{i\to j}^{(v)}$ by a function $\psi_{M}$ which is a chain of two GVPs:
$m_{i\to j}^{(s)},m_{i\to j}^{(v)}=\psi_{M}\left(\left[s_{i}^{(l)}:e_{ij}^{(l)}:d_{ij}^{(l)}\right],\left[v_{i}:\frac{x_{i}^{(l)}-x_{j}^{(l)}}{d_{ij}^{(l)}}\right]\right)$ (13)
Where $d^{(l)}_{ij}$ is the distance between nodes $i$ and $j$ at update block $l$. In the FlowMol implementation, the distance $d_{ij}$ is replaced with a radial basis distance embedding before being fed into the GVPs or MLPs. Following the message generation step, an update of node scalar and vector features can be accomplished by aggregating generated messages:
$s_{i}^{(l+1)},v_{i}^{(l+1)}=\mathrm{LN}\left([s_{i}^{(l)},v_{i}^{(l)}]+\psi_{N}\left(\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}\right)\left[m_{j\to i}^{(s)},m_{j\to i}^{(v)}\right]\right)$ (14)
Where $LN$ stands for LayerNormalization operation, $\psi_{N}$ is a chain of three GVPs.
Node position update.
NPU block updates node positions by taking node-wise operations on the updated node scalar and vector features:
$x_{i}^{(l+1)}=x_{i}^{(l)}+\psi_{P}\left(s_{i}^{(l+1)},v_{i}^{(l+1)}\right)$ (15)
Where $\psi_{P}$ is a chain of three GVPs in which the final output has 1 vector and 0 scalar features.
Edge feature update.
Edge features are updated by edge-wise operations that take the updated node scalar features and node distance as the inputs:
$e_{ij}^{(l+1)}=\mathrm{LN}\left(e_{ij}^{(l)}+\mathrm{MLP}\left(s_{i}^{(l+1)},s_{j}^{(l+1)},d_{ij}^{(l+1)}\right)\right)$ (16)
Models are trained using the settings detailed in the ‘Hyperparameters and training settings’ section.
Property embedding operations
@@
Conditional generation can be framed as sampling a property-conditional target distribution $p_{1}(x|k)$ over molecular graphs, where $k$ is a property. Previous approaches to conditional generation concatenate a property value to the node features of a molecular graph, training the model jointly on the property and the graph [hoogeboom2022, xu2023, song2023]. As a result, the learned denoising process is conditioned on the property. This approach requires modifying the input dimension of the denoiser to include the additional property value, which is subsequently discarded in the final output. However, directly appending the property to the node features may oversimplify the interaction between property values and molecular graphs. As shown in Figure 1, our implementation maps a scalar property, $k$, to a high-dimensional property embedding space $P=\phi_{\mathrm{prop}}(k)$ using a MLP, $\phi_{\mathrm{prop}}$. The dimensionality of the property embedding matches that of the node scalar features, allowing seamless integration without requiring additional dimensional matching. To encode interactions between the property embedding and node scalar features in molecualr graphs, we propose five distinct operations, including ‘Concatenation’, ‘Sum’, ‘Multiply’, ‘Concatenate + Sum’, and ‘Concatenate + Multiply’. Depending on the operation, if the combined feature vector $[A,C]$ maintains its original dimensionality after interacting with the property embedding, no further processing is applied; otherwise, an MLP is used to project it back to the original dimensionality. Since the property embedding interacts exclusively with node scalar features, leaving position vectors unchanged, the resulting property-conditioned graph retains the SE(3) equivariance. Moreover, the loss function preserves its original form, and the trained SE(3) GNN defines a valid conditional generative process.
@@
Specifically, in the ‘Sum’ operation, the property embedding is simply added to the node scalar features. The ‘Multiply’ operation performs an element-wise Hadamard product $\odot$’. To ensure the multiplicative factors remain within a bounded range, the property embedding is passed through a sigmoid function and then shifted by 0.5, resulting in multipliers in the interval $[0.5,1.5]$. Both ‘Sum’ and ‘Multiply’ preserve the original feature dimensionality. In contrast, operations involving ‘Concatenate’ increase the dimensionality of the node scalar features, and an MLP transformation $\varphi_{\theta}$ is required to project the combined features back to the original dimension. For the composite operations, ‘Concatenate + Sum’ and ‘Concatenate + Multiply’, the property embedding is first concatenated to the node features, followed by the respective arithmetic operation. In both cases, the final output has the desired dimensionality, either by design or via the subsequent MLP transformation.
Gaussian expansion
@@
We incorporate a fixed, non-trainable Gaussian expansion layer to enrich the property representations prior to being mapped to a property embedding via an MLP $\phi_{\mathrm{GE}}$ [gebauer2022]. The Gaussian expansion maps a scalar property value into a fixed-length vector using a set of Gaussian basis functions. Given a property $k$ with value $\tau_{k}$, the expanded representation takes the form:
@@
Figure 1d shows an example of a Gaussian expansion for the HOMO-LUMO gap. In this example, five Gaussian basis functions ($n_{g}=5$) are used, with centers evenly spaced between the minimum and maximum gap values in the QM9 dataset: $\tau_{\mathrm{min}}=0.0246$ and $\tau_{\mathrm{max}}=0.6221$ Hartree. The width of each Gaussian is set by the spacing parameter $d=(\tau_{\mathrm{max}}-\tau_{\mathrm{min}})/(n_{g}-1)$. The top panel of Figure 1d shows the five Gaussian basis functions, with their centers marked by vertical dashed gray lines. In the bottom panel, we pick two example property values, namely 0.1, and 0.5 Hartree, and show the expanded feature vectors as bar plots indicating the activation of each Gaussian basis function. For instance, if the gap is 0.1 Hartree, the first two Gaussian functions (centered closest to 0.1) have higher responses than the remaining functions, whose centers are farther away. Similar activation patterns are observed for the other two property values.
Inference
ID tasks.
To generate paired samples of property values and numbers of atoms, we employed a two-stage sampling procedure. First, atom counts were sampled according to their empirical distribution in the training dataset. Specifically, the frequency of each unique atom count was computed and used to construct a categorical probability distribution, from which atom counts were drawn.
Conditional on a given atom count, the associated property values were sampled using a histogram-based approximation of their empirical distribution. For each unique atom count, the observed property values were discretized into a fixed number of bins like 1000. The counts within each bin were normalized to form a categorical distribution. During sampling, a bin index was first drawn from this categorical distribution, after which a continuous property value was assigned by uniformly sampling within the range of the selected bin.
This approach yields paired samples of atom counts and property values that respect both the marginal distribution of atom counts and the conditional distribution of properties given the atom count. To make a fair comparison across different property embedding methods, the same sampled target values and atom counts were used.
OOD tasks.
The property values in the OOD tasks are chosen as the 99th quantile of the QM9 training data and the numbers of atoms are selected for the property ranges bounded by the 97th quantile and the maximum (100th quatile) property value. During inference, PropMolFlow models use Euler’s method with 100 evenly spaced time steps to integrate the learned velocity field.
Datasets, baselines and evaluation metrics
Datasets
@@
We trained PropMolFlow on the QM9 data with explicit hydrogen atoms [ramakrishnan2014quantum, wu2018moleculenet]. We did not evaluate PropMolFlow on the GEOM-Drugs dataset, which is commonly used for unconditional generation and includes larger molecules, because it only provides quantum-mechanical energy values and lacks other relevant properties [axelrod2022]. The original QM9 data contains 133,885 molecules, and each molecule contains up to 9 heavy atoms and consists of 3–29 atoms (average: 18 atoms), composed of up to 5 elements: C, N, O, F, and H. Molecules are fully optimized and molecular properties are obtained with DFT at the level of B3LYP/6-31G(2df,p) [becke1993, lyp1988]. For property-guided generation, we considered six molecule properties: polarizability ($\alpha$), HOMO-LUMO gap, HOMO energy, LUMO energy, dipole moment ($\mu$), and heat capacity ($C_{v}$). Precise definitions of each property can be found in Supplementary Section LABEL:sec:more_details. Upon inspection, we discovered numerous inconsistencies: invalid bond orders and non-zero net charges in roughly 30,000 molecules, despite QM9’s requirement for charge-neutral, closed-shell valency [ramakrishnan2014quantum]. We corrected these discrepancies by reassigning bond orders or charges to enforce valency-charge consistency for the vast majority of entries and used the corrected data to train our PropMolFlow models. After the correction, we improved the structural validity of generated molecules substantially (Supplementary Table LABEL:tab:data_and_metric_compare). Previous works that used bond orders overlooked this issue [vignac2023, dunn2024-2]; instead, they modified evaluation metrics for models trained on the problematic SDF data against the one used by Hoogeboom et al. [hoogeboom2022] to tolerate chemically inconsistent bond–charge combinations, hence inflating molecule stability estimates contrary to QM9’s intended constraints. A detailed comparison between models using previous problematic data and the corrected data, together with the comparison between previous stability metrics and the revised ones, can be found in Supplementary Table LABEL:tab:data_and_metric_compare. A detailed description of our procedure and the alignment with the original QM9 XYZ data is provided in Supplemental Section LABEL:sec:qm9_data_fix. The revised QM9 SDF data is provided on Zenodo [zenodo_propmolflow]. We used RDKit’s sanitization function [rdkit] to filter chemically valid molecules on the QM9 SDF file with bond and charge fix, resulting in a curated set of 133k molecules. After random shuffling, we split the dataset into 100k training, 20k validation, and 13k test samples. PropMolFlow models and GVP property predictors are trained on disjoint 50k and 50k datasets from the 100k training data.
Baselines
We compare PropMolFlow against several recent property-guided molecular generation methods based on diffusion models and flow matching, including EEGSDE, GCDM, GeoLDM, JODO, and EquiFM [bao2023, morehead2024, xu2023, huang2024, song2023]. EEGSDE (Equivariant Energy Guided Stochastic Differential Equations) use an energy function to guide the diffusion for property-guided generation [bao2023]. EquiFM is the first equivariant flow matching model for 3D molecule generation [song2023]. GeoLDM represents the first latent diffusion model for molecular geometry generation [xu2023]. GCDM constructs an SE(3) equivariant geometry-complete diffusion model using geometry-complete perceptrons for molecule generation [morehead2024, moret2023]. JODO uses a diffusion graph transformer for joint generation of 2D molecular graphs and 3D geometries. EEGSDE, EquiFM, GeoLDM, and GCDM do not include bond orders in their molecular graphs, while JODO incorporates both bond existence and bond orders. All five baseline models omit atomic charges in their conditional generation, in contrast to PropMolFlow, which can generate molecules with non-zero atomic charges.